OpenStudy (anonymous):

A rectangle is three times as long as it is wide. If the length is decreased by 2 cm and the width is increased by 3 cm, the new rectangle formed has an area of 42 cm^2. What are the dimensions of the original rectangle?

OpenStudy (dbzfan836):

Hmmm... have you tried finding out the answer by using the trial-error-success method?

OpenStudy (dbzfan836):

you have to try to guess the dimensions, then test it to get the area.

OpenStudy (anonymous):

So I plugged in 9 as the length, and 3 as the width. 9 by 3 is 27, but if you subtract 2 from the length (9-2 = 7) and add 3 to the width (3+3=6) and multiply those answers together: 6 x 7 = 42. Then you would cube that, making it 42 cm cubed. Correct?

OpenStudy (dbzfan836):

ok... I'm gonna use the trial method... 42 / 7 [7 is the length] = 6 [6 is the width] 7+2=9 [9 is the previous length] 6-3=3 [3 is the previous width]

OpenStudy (dbzfan836):

No, that would be squared.

OpenStudy (anonymous):

Right, I meant squared. Sorry!

OpenStudy (dbzfan836):

Yes, that is correct, good job!

OpenStudy (dbzfan836):

you have found the answers. here they are simple and sweet: L=9 W=3